Find the solutions to the equation: x+7/x=8. Answer is either: A.-8,1 B.1,7 C.2,4 D.3,4. Please advise, thank you!
Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

x+7/x=8 or (x~2+7)/x=8 or x~2+7-8x=0 or x~2-8x+7=0 or (x-7)(x-1)=0 or x=1 & x=7 ans. B)1,7

Answer by stanbon(48568)

x+(7/x)=8
Multiply through by x to get:
x^2+7=8x
x^2-8x+7=0
(x-7)(x-1)=0
x=7 or x=1
Cheers,
Stan H.